OpenStudy (anonymous):

How do i solve for x in..... 8^3x-9=4^10x-7

OpenStudy (anonymous):

isolate x on one side and put everything else on other side.......

OpenStudy (anonymous):

8^3x=4^10x+2

OpenStudy (anonymous):

512x=10000x+2 -9488x=2 x=-2/9488

OpenStudy (anonymous):

3x-9 and 10x-7 are exponents of 8 and 4 respectively.. my answer book is saying the answer is wrong so im just trying to find the problem

OpenStudy (anonymous):

ohh...okay sorry

OpenStudy (anonymous):

do you have another answer?

OpenStudy (anonymous):

nope sorry

hartnn (hartnn):

is it \(\huge 8^{3x-9}=4^{10x-7}\) ?

hartnn (hartnn):

if yes, you can write that as \(\huge 2^{3(3x-9)}=2^{2(10x-7)}\) then equate the exponents to get 3(3x-9)=2(10x-7) solve for x now.